A and B have a combined age of 41 but A is two more than twice B's age...what are their ages? the formula is A+B=41,A is two more than twice B's age..
A = 2b + 2

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

A and B have a combined age of 41 but A is two more than twice B's age...what are their ages? the formula is A+B=41,A is two more than twice B's age..
A = 2b + 2
--------------------------
A + B = 41
A -2B = 2
----
Subtract and solve for B:
3b = 39
B = 13
------------
Solve for A:
A + B = 41
A + 13 = 41
A = 28
-------------------------
Cheers,
stan H.
=====

Answer by solver91311(24713)



You have both equations set up correctly, now just substitute:





So:



goes in place of in the first equation resulting in:



Solve for . Then calculate . Check your work by calculating which, given proper work, should equal

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism